OpenStudy (anonymous):

A rope is cut into three pieces marked A, B, and C. The length of the pieces are in the ratio 4 : 6 : 8. The rope is 36 feet long. What is the length of A? 4 ft 8 ft 16 ft 22 ft

OpenStudy (anonymous):

last one

OpenStudy (dayakar):

ratio of three pieces = 4:6:8 (given)

OpenStudy (dayakar):

let lengh of A = 4x lengh of B =6x lengh of C 8x

OpenStudy (dayakar):

total lengh =36 ft (given)

OpenStudy (dayakar):

4x+6x+8x = 36

OpenStudy (jango_in_dtown):

so x=2 gives 4x=8

OpenStudy (dayakar):

18x =36 x=36/18 x=2

OpenStudy (dayakar):

A=4x a= 4*2 A=8ft

OpenStudy (dayakar):

is it clear